# Stable Diffusion Segmentation for Biomedical Images with Single-step Reverse Process

### v1

@@ 3.3.3 Stability Evaluation
Table. 5 showcases SDSeg’s significant stability across these tests, underscoring its robust generalization ability despite different initial noises.
@@ 4 Conclusion
Supplemental Materials
Figure 4: Illustration of our Stability Evaluation on REF. To prepare for the evaluation, we first conduct $M$ times of inference process. Then, Dataset-level Stability (DS)(in red) is evaluated on every two sets of the inference results; Instance-level Stability (IS)(in blue) is estimated on every two segmentation maps of each image conditioning. In practice, we choose the first inference results to conduct the DS evaluation with the other $M-1$ sets, and the first instance of segmentation result to conduct IS evaluation with the other $M-1$ segmentation maps.
Figure 5: Visualization of the predicted probability maps in reverse process on STS. The horizontal axis denotes DDIM sampling steps. Upper two rows: SDSeg trained for only 5,000 steps, sub-optimal results appear in early reverse process; Lower two rows: SDSeg trained for 50,000 steps, DDIM sampler generates fine and stable results during the entire reverse process. This demonstrates that SDSeg can generate great results under limited steps of the reverse process.

### v3

@@ 3.3.3 Stability Evaluation
Table. 5 showcases SDSeg’s significant stability across these tests, underscoring its reliability in segmentation despite different initial noises.
@@ 4 Conclusion
{credits}
4.0.1 Acknowledgements
This study was funded by the Program of Science and Technology of Guangdong (Grant No. 2020B1111170009), the 2022 Industrial Technology Basic Public Service Platform Project of China (Grant No. 2022-228-219), and the Fundamental Research Funds for the Central Universities, Sun Yat-sen University (Grant No. 23xkjc016).
[…]
The authors have no competing interests to declare that are relevant to the content of this article.
Appendix 0.A Stability Evaluation
Figure 4: Illustration of our Stability Evaluation on REF. We first conduct $M$ times of inference process to prepare for the evaluation. Then, Dataset-level Stability is evaluated on every two sets of the inference results; Instance-level Stability is estimated on every two segmentation maps of each image conditioning.
Appendix 0.B Qualitative Analysis
Figure 5: From top to bottom: Visualization of the predicted probability maps in reverse process on CVC, BTCV, and KSEG (SDSeg trained for 50,000 steps). The horizontal axis denotes DDIM sampling steps. DDIM sampler generates fine and stable results during the entire reverse process. This demonstrates that SDSeg can generate great results under limited steps of the reverse process.
@@ Appendix 0.B Qualitative Analysis
Appendix 0.C The architecture of the trainable vision encoder
We use a KL-regularized autoencoder model with the downsampling rate $r=\frac{H}{h}=\frac{W}{w}=8$. The proposed trainable vision encoder has the same network architecture as the autoencoder model’s encoder. Specifically, the trainable vision encoder’s architecture can be separated into three blocks: the Downsampling block (Table. 9), the Mid block (9) and the Out block (9).
In Table. 9, ‘Conv 3$\times$3’ denotes convolution block with kernel size 3, ‘ResBlock’ represents the building block in ResNet, and ‘Down’ corresponds to downsampling. In Table. 9, ‘Attention’ denotes self-attention block.
Table 7: The architecture of the Downsampling block.
input $\mathbb{R}^{H\times W\times 3}$
Conv 3$\times$3 $\mathbb{R}^{H\times W\times C}$
ResBlock×$\times 2$+Down $\mathbb{R}^{\frac{H}{2}\times\frac{W}{2}\times C}$
ResBlock×$\times 2$+Down $\mathbb{R}^{\frac{H}{4}\times\frac{W}{4}\times 2C}$
ResBlock×$\times 2$+Down $\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 4C}$
ResBlock×$\times 2$ $\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 4C}$
Table 8: The architecture of the Mid block.
input $\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 4C}$
ResBlock $\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 4C}$
Attention $\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 4C}$
ResBlock $\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 4C}$
Table 9: The architecture of the Out block.
input $\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 4C}$
GroupNorm $\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 4C}$
Conv 3$\times$3 $\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 2Z}$
Conv 1$\times$1 $\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times Z}$
The input segmentation map $X\in\mathbb{R}^{H\times W\times 3}$ successively goes through these three blocks to get its corresponding latent representation $z\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times Z}$, where $C=128$ is the channel dimension of the vision encoder, and $Z=4$ is the channel dimension of the latent representation.